# Analysis and Validation of Image Search Engines in Histopathology

### v1

@@ Abstract
Searching for similar images in archives of histology and histopathology images is a crucial task that may aid in patient matching for various purposes, ranging from triaging and diagnosis to prognosis and prediction. Whole slide images (WSIs) are highly detailed digital representations of tissue specimens mounted on glass slides. Matching WSI to WSI can serve as the critical method for patient matching. In this paper, we report extensive analysis and validation of four search methods bag of visual words (BoVW), Yottixel, SISH, RetCCL, and some of their potential variants. We analyze their algorithms and structures and assess their performance. For this evaluation, we utilized four internal datasets ($1269$ patients) and three public datasets ($1207$ patients), totaling more than $200,000$ patches from $38$ different classes/subtypes across five primary sites. Certain search engines, for example, BoVW, exhibit notable efficiency and speed but suffer from low accuracy. Conversely, search engines like Yottixel demonstrate efficiency and speed, providing moderately accurate results. Recent proposals, including SISH, display inefficiency and yield inconsistent outcomes, while alternatives like RetCCL prove inadequate in both accuracy and efficiency. Further research is imperative to address the dual aspects of accuracy and minimal storage requirements in histopathological image search.
1 Introduction
@@
With the surge of development in digital pathology over recent years, there has been an increase in demand for image search solutions to allow pathologists to quickly compare tissue images and learn from previously diagnosed and treated cases. This process is expected to resemble peer review and consultation to acquire second opinions. However, there are several challenges and roadblocks in image search in digital pathology archives that make creating an efficient and semantically accurate image search engine extremely difficult. The first issue is the extremely large nature of whole slide images (WSIs), which are often gigapixel files. So far, the ideal search engines must offer a method to “divide” or break up these WSIs into smaller more manageable “patches” or tiles. That implies that an ideal search engine must use a form of unsupervised algorithm to divide each WSI into a small set of patches. That only a few patches have to be selected to represent the WSI is a computational and storage requirement that would directly impact the adoption of search technologies in hospitals. Subsequently, to index the WSI, the selected small set of patches are then often run through a pre-trained, or fine-tuned deep network that extracts feature vectors, also called embeddings or deep features, from each patch. This brings us to the second issue regardless of the quality of the deep features. The ideal search engine should also offer an encoding scheme that allows for efficient storage and fast comparison of the deep features. Efficiency is not solely determined by speed but also by the amount of memory required to index images. The latter aspect has often been disregarded in literature, favouring fast search.
Different image search methods have been proposed in recent years. Some of them are complete search pipelines, whereas the others rather just some components of a search engine.
@@
In this paper, we report the extensive analysis and validation of four frameworks and some of their variations for search in histopathology archives. We examine the bag of visual words (BoVW) [1], Yottixel [2], Self-Supervised Image Search for Histology (SISH) [3], and Retrieval with Clustering-guided Contrastive Learning (RetCCL) [4], in several configurations using four private and three public datasets. We first provide a brief overview of content-based image retrieval (CBIR) literature to establish the historical perspective. We then motivate the selection of these four search frameworks and explain how they work. We then report a comprehensive set of results to illuminate the algorithmic structure and efficiency and accuracy performance of each approach. This is the first rigorous and comparative analysis and validation of the most relevant image search frameworks in histopathology. We expect that the results reported in this paper will make clear where we stand in image search as a community and will help pathologists and other stakeholders in making decisions about which ones, if any, are intrinsically more tailored to clinical workflow and worth testing for research purposes.
2 Image Search in Histopathology
@@
The field of information retrieval, particularly in the medical context, is a well-studied field and has undergone significant advancements [5, 6, 7, 8, 9, 10, 11]. In recent years, Content-Based Image Retrieval (CBIR) in general and image search techniques have emerged as potentially indispensable tools in the field of histology and histopathology in particular, both for traditional light microscopy and digital pathology [12, 2, 3, 1, 13, 4, 14, 15]. Well-designed and efficient image search technologies may enable pathologists to effectively access, analyze, and compare large volumes of tissue images, enhancing the accuracy of diagnoses and treatment planning [16, 17, 18].
@@
A thorough validation of image search solutions in histopathology in recent years seems to be missing. Looking at the vast and rich literature on CBIR, one may be inclined to include many methods in a validation study. However, considering specific requirements of medical CBIR decreases the number of candidate image search solutions drastically. Bridging from general medical image analysis to digital pathology leaves us with a handful of candidates. We established only one condition to make the final selection: the image search candidates in histopathology worth examination should introduce, or at least employ a distinct “divide” strategy to be feasible in clinical settings; dividing a WSI into patches and selecting a small set of those patches is the indispensable prerequisite for patient matching. Pertinent to the validation process, it is of secondary importance whether a search strategy has introduced its divide approach or borrows the divide from literature. Based on all these considerations, we selected four image search frameworks, namely BoVW [1], Yottixel [2], SISH [3] and RetCCL [4]. As a recent example, we did not include SMILY [13] in our validation because it cannot process WSIs due to the lack of a divide stage. SMILY assumes all WSI patches should be processed, a task that requires prohibitively extensive resources as most WSIs consist of several hundred if not several thousand patches.
2.1 BoVW
@@ 2.4 RetCCL
3 Results
@@
We compared the four selected search approaches in terms of their algorithmic structure, search capabilities, training, and test data used to validate them. More importantly, we conducted extensive experiments to compare their search performance. As for algorithmic structure comparison, we compared the patching methods they used to break down the WSI and perform indexing and matching. Overall, by comparing their search capabilities, we determined whether they are able to process/search for both patch and WSI matching. Training and testing data comparisons encompass a detailed inventory of the data each search engine was trained with, or validated on. Finally, we used both high-quality clinical datasets and well-known public datasets to quantify the search performance of each method. We examined accuracy, time measurements for indexing and search, robustness, and storage requirements as main performance benchmarking.
3.1 Algorithmic Structure
@@
Table 1 offers an overview of the algorithmic structures employed by the four search frameworks (also see Figures 1,2,3 and 4 that visualize only the first three columns of Table 1). Each approach is detailed across six key stages: Divide, Features, Encoding, Space, Matching, Speed, and Post-Processing. To analyze the algorithmic structure, we considered both the descriptions in the corresponding paper, as well as details of their implementation. We did not list “pre-processing” as all four methods start with segmenting the slide into tissue and non-tissue regions.
Table 1: Algorithmic comparisons of all search methods (also see Figures 1,2, 3 and 4)
@@
Divide – Since using all WSI patches for the downstream tasks is inefficient in terms of both computation and storage (as this is the case for SMILY [13]), search methods must employ techniques to select a subset of the patches, also known as the “divide” stage in a divide & conquer attempt to master the gigapixel nature of WSIs. While BoVW uses a random sampling approach to select visual words (which are much smaller than patches), Yottixel uses its “mosaic” approach, an unsupervised approach to get a small set of patches. SISH does not offer any new solution and uses Yottixel’s mosaic, the same as RetCCL. Yottixel performs color and spatial clustering to select representative patches, whereas RetCCL uses deep feature and spatial clustering, a much more expensive variation.
Feature Extraction – Every search method has to utilize deep neural networks to obtain deep features from the image data, i.e., patches. RetCCL uses its clustering-guided contrastive learning (CCL) trained network using TCGA images to extract features for all patches prior to the mosaic generation step since it utilizes them for the divide step. In contrast, Yottixel and SISH both generate feature vectors only for the mosaic patches. Yottixel uses a pre-trained DenseNet, which can be swapped by any other network. SISH also extracts features using DenseNet, with the addition of an autoencoder trained on TCGA images. The rationale behind integrating the autoencoder alongside DenseNet in SISH is not sufficiently justified. It has been argued that SISH does not use any self-supervision for training the autoencoder, hence the paper may have introduced a misnomer [30]. Originally, BoVW used local binary patterns (LBP) as features but for the sake of comparison, we replaced it with an autoencoder so all the search engines utilize deep features.
@@
Post-Processing – Generally, the search process concludes once results have been retrieved and sorted based on their similarity to the query. However, both SISH and RetCCL utilize an additional ranking algorithm after the search, which eliminates patches deemed less likely to be similar to the query. The inclusion of ranking as a post-processing step is evidently motivated by a lack of confidence in the search results. This approach seems misguided, as relying on post-search ranking implies a lack of trust in the initial results. The authors of SISH stress the significance of ranking for their outcomes, indicating that incorporating an autoencoder into the Yottixel pipeline has not yielded any discernible benefits, despite the paper emphasizing the autoencoder’s role.
3.2 Search Capabilities
@@
It is an indispensable capability of a search engine in histopathology to be able to perform both patch search and WSI search. Clearly, WSI-2-WSI matching (WSI search) is a more difficult task compared to patch-to-patch matching. Also, WSI search is pivotal for patient matching. Both BoVW and Yottixel stand out as comprehensive solutions with the ability to perform both patch and WSI searches. In contrast, SISH and RetCCL, while being able to process and search patches due to using the Yottixel’s mosaic, lack the capability to perform WSI-2-WSI matching. This in fact stems from their dependence on ranking patches after the search to find the label for the query WSI. Therefore, although the label found by SISH and RetCCL may be restricted to top-$n$ patches, this does not constitute true WSI-2-WSI matching. In the case of SISH, even without post-search ranking, the tree used for fast processing cannot be restricted to top-$n$ WSIs.
In summary, BoVW creates a histogram for an image (either a patch or a WSI). The histogram can be calculated for both patch and WSI (Figure 5, left). Hence, BoVW can process and match both patches and WSIs. Yottixel creates a mosaic of patches for each WSI and uses the “median of minimum distances” to get a single distance metric for WSI comparison (Figure 5, middle); Hence, Yottixel can match both patches and WSIs. Both SISH and RetCCL use Yottixel’s mosaic but their ranking algorithms essentially force them to perform patch matching although the ranking may be configured the select one patch per WSI to emulate WSI selection. They take a query patch and compare it with every other patch in every other WSI (Figures 5, right). SISH pre-raking tree processing is not WSI-conform either. Hence, SISH and RetCCL are rather patch classifiers when using post-search ranking, hence not usable for patient matching in a clinical setting.
@@
3.3 Training/Test Data Comparison
The data used for any type of training, the test data, and the user feedback experiments are all important factors in evaluating search engines. Table 2 offers an overview of the data utilized by the four search engines. BoVW uses Kimia Path24111https://github.com/RhazesLab/kimia_path24 as testing data, comprising 24 different anatomic sites [34]. Notably, BoVW, as reported in [1], does not use any deep features and does not conduct user tests. Yottixel uses a pre-trained model, namely DenseNet. Hence, it has been validated on entire TCGA images, both diagnostic slides (FFPE slides) and frozen sections, in addition to some private data from UPMC, accompanied by some user tests involving three pathologists. The data Yottixle has been validated on covers a wide range of anatomic sites, totaling 27 primary sites, and an extensive number of subtypes. It is notable that since the pre-trained DenseNet121 has been trained on ImageNet, which is a non-pathological dataset, all Yottixel tests can be considered external validation. SISH draws its training data from TCGA and conducts testing with a diverse dataset comprising TCGA, CPTAC, and BWH data, without involving user tests. It focuses on a reduced set of anatomic sites (13) but accounts for 56 different subtypes. It is not clear how the training and testing TCGA data for SISH have been separated. RetCCL uses TCGA and PAIP data for training and only TCGA data for testing, without user tests. RetCCL considers 13 anatomic sites but focuses on a more limited set of subtypes.
Table 2: Training and Test Data Comparison
Training Data Test Data User Test

### v2

@@ Abstract
Searching for similar images in archives of histology and histopathology images is a crucial task that may aid in patient tissue comparison for various purposes, ranging from triaging and diagnosis to prognosis and prediction. Whole slide images (WSIs) are highly detailed digital representations of tissue specimens mounted on glass slides. Matching WSI to WSI can serve as the critical method for patient tissue comparison. In this paper, we report extensive analysis and validation of four search methods bag of visual words (BoVW), Yottixel, SISH, RetCCL, and some of their potential variants. We analyze their algorithms and structures and assess their performance. For this evaluation, we utilized four internal datasets ($1269$ patients) and three public datasets ($1207$ patients), totaling more than $200,000$ patches from $38$ different classes/subtypes across five primary sites. Certain search engines, for example, BoVW, exhibit notable efficiency and speed but suffer from low accuracy. Conversely, search engines like Yottixel demonstrate efficiency and speed, providing moderately accurate results. Recent proposals, including SISH, display inefficiency and yield inconsistent outcomes, while alternatives like RetCCL prove inadequate in both accuracy and efficiency. Further research is imperative to address the dual aspects of accuracy and minimal storage requirements in histopathological image search.
Index Terms:
Image Search, Image Retrieval, Deep Learning, Histopathology, Histology
I Introduction
@@
With the surge of development in digital pathology over recent years, there has been an increase in demand for image search solutions to allow pathologists to quickly compare tissue images and learn from previously diagnosed and treated cases. This process is expected to resemble peer review and consultation to acquire second opinions. However, there are several challenges and roadblocks in image search in digital pathology archives that make creating an efficient and semantically accurate image search engine extremely difficult. The first issue is the extremely large nature of whole slide images (WSIs), which are often gigapixel files. So far, the ideal search engines must offer a method to “divide” or break up these WSIs into smaller more manageable “patches” or tiles. That implies that an ideal search engine must use a form of unsupervised algorithm to divide each WSI into a small set of patches. That only a few patches have to be selected to represent the WSI is a computational and storage requirement that would directly impact the adoption of search technologies in hospitals. Subsequently, to index the WSI, the selected small set of patches are then often run through a pre-trained, or fine-tuned deep network that extracts feature vectors, also called embeddings or deep features, from each patch. This brings us to the second issue regardless of the quality of the deep features. The ideal search engine should also offer an encoding scheme that allows for efficient storage and fast comparison of the deep features. Efficiency is not solely determined by speed but also by the amount of memory required to index images. The latter aspect has often been disregarded in literature, favouring fast search algorithms.
@@
Different image search methods have been proposed in recent years. Some of them are complete search pipelines, whereas the others are rather just some components of a search engine. In this paper, we report the extensive analysis and validation of four frameworks and some of their variations for search in histopathology archives. We examine the bag of visual words (BoVW) [1], Yottixel [2], Self-Supervised Image Search for Histology (SISH) [3], and Retrieval with Clustering-guided Contrastive Learning (RetCCL) [4], in several configurations using four private and three public datasets. We first provide a brief overview of content-based image retrieval (CBIR) literature to establish the historical perspective. We then motivate the selection of these four search frameworks and explain how they work. We then report a comprehensive set of results to illuminate the algorithmic structure and efficiency and accuracy performance of each approach. This is the first rigorous and comparative analysis and validation of the most relevant image search frameworks in histopathology. We expect that the results reported in this paper will make clear where we stand in image search as a community and will help pathologists and other stakeholders in making decisions about which ones, if any, are intrinsically more tailored to clinical workflow and worth testing for research purposes.
II Image Search in Histopathology
@@
The field of information retrieval, particularly in the medical context, is a well-studied field and has undergone significant advancements [5, 6, 7, 8, 9, 10, 11, 12, 13]. In recent years, Content-Based Image Retrieval (CBIR) in general and image search techniques have emerged as potentially indispensable tools in the field of histology and histopathology in particular, both for traditional light microscopy and digital pathology [14, 2, 3, 1, 15, 4, 16, 17]. CBIR searches and retrieves images from a database based on their visual features. In contrast to using textual metadata like captions, keywords, or tags, CBIR systems analyze the visual content of images (i.e., the pixels) to identify objects’ similarities and differences. Generally, various features are extracted from images (e.g., color, texture, shapes) to index and organize the images in the database [18, 19, 20]. In this work, we mainly use “image search” as equivalent to CBIR.
CBIR applications for pathology have been investigated even before the emergence of digital pathology and deep learning [21, 22, 23]. In those techniques, generally handcrafted features would be used to characterize pathology images [24, 25]. Computer vision research provided a wealth of handcrafted features, such as SIFT [26], SURF [27], and LBP [28]. It is accepted empirically that deep features are superior to handcrafted features [29, 30, 31].
@@
Well-designed and efficient image search technologies may enable pathologists to effectively access, analyze, and compare large volumes of tissue images, enhancing the accuracy of diagnoses and treatment planning [32, 33, 34]. A thorough validation of image search solutions in histopathology in recent years seems to be missing. Looking at the vast and rich literature on CBIR, one may be inclined to include many methods in a validation study. However, considering specific requirements of medical CBIR decreases the number of candidate image search solutions drastically. Bridging from general medical image analysis to digital pathology leaves us with a handful of candidates. We established only one condition to make the final selection: the image search candidates in histopathology worth examination should introduce, or at least employ a distinct “divide” strategy to be feasible in clinical settings; dividing a WSI into patches and selecting a small set of those patches is the indispensable prerequisite for patient tissue comparison (one may also represent the entire WSI with one vector [35, 36, 37]). Pertinent to the validation process, it is of secondary importance whether a search strategy has introduced its own divide approach or borrows the divide from literature. Based on all these considerations, we selected four image search frameworks, namely BoVW [1], Yottixel [2], SISH [3] and RetCCL [4].
As recent examples, we did not include SMILY [15], HSDH (histopathology Siamese deep hashing) [38] and High-Order Correlation-Guided SelfSupervised Hashing-Encoding Retrieval (HSHR) [39] in our validation. Table I provides an overview of the challenges with these methods. The main practical reason for not evaluating these three methods is that their code and/or trained models are not publicly available such that we cannot reliability reproduce their results. As well, they lack a divide stage and follow a brute force approach to WSI representation. HSDH is mainly focused on a Siamese network to process single patches. SMILY and HSHR assume all WSI patches should be processed, a task that requires prohibitively extensive resources as most WSIs consist of several hundred if not several thousand patches (see Section III-A where the relationship of storage and AI democratization is reviewed).
TABLE I: Excluded Search Methods. The ‘brute force’ approach does not follow a divide & conquer scheme but processes all patches of a WSI.
Methods Patching Code Model(s)
SMILY [15] Brute force not available not available
HSDH [38] Brute force not available not available
HSHR [39] Brute force available not available
II-A BoVW
@@ II-D RetCCL
III Methods
This section describes the process of conducting our experiments, how the search engines were implemented, what pre-trained networks were involved, and what datasets were used.
III-A Experimental Setup
Search accuracy was measured among the top-1, majority vote among top-3, and majority vote among top-5 WSI matching results. The $F_{1}$ score was used as a measure to quantify both precision and recall. Additionally, the macro-averaged accuracy ($\bar{F}_{1}$) was calculated for each search engine in each experiment. All the experiments regarding SISH, Yottixel, and RetCCL were carried out using a leave-one-patient-out approach which is the extreme form of cross-validation where the number of folds is equal to the number of patients in the dataset. Since the BoVW approach requires training an autoencoder for each dataset, the experiments for this specific search engine were carried out via a 5-fold cross-validation. We also measured the speed of each search engine in indexing and searching, separately. The average time taken to create the index was reported in minutes and the average search times (including both indexing and matching time for the query) were reported in seconds. To quantify reliability, we also counted how many times each search engine failed to process a WSI in each dataset. The failed WSIs were excluded from that specific experiment of search engine and dataset. We also measured the storage need of each search engine, separately. After indexing all the WSIs in each dataset, we measured the total disk space each search engine occupied to index the WSIs, excluding failed ones, to obtain the average disk space needed for each WSI. We also extrapolated the estimated amount of disk space needed to index one million WSIs using the estimated average. The significance of storage is amplified by the desire for democratization of AI and the research community ambitions for inclusion and equity. WSIs being much larger than radiology images have to be handled on top of “already financially and operationally stressed healthcare system” [52]. Most labs in developing countries do not have resources to store WSIs [53]. Asking them for excessive extra storage to run a search engine appears illusory, a challenge that is widely recognized [54].
III-B Implementation
We cannot claim to have trained the best autoencoder for BoVW. However, by introducing deep features into BoVW, we were able to include this well-known framework in the validation process.
Yottixel. We used the sample Python code publicly published on KimaLab’s website [^1] https://kimialab.uwaterloo.ca/kimia/index.php/yottixel/. The published code demonstrates the steps needed to index and search one sample WSI. We used the exact code to index all the WSIs iteratively. We then utilized the saved indexed images to search for a query WSI. The sample code uses a pre-trained DenseNet121 for feature extraction which is the original proposed version of Yottixel. As previously explained in the experiment setup section, we also tested two new variants of the original Yottixel. One variant was created by swapping the DenseNet121 with KimiaNet as the feature extractor. The other variant was implemented to use post-search ranking, as it is already implemented in SISH. To do so, we got the patch-level features extracted by Yottixel and applied the same ranking used by SISH. It is noteworthy that this change rendered Yottixel unable to match WSI-to-WSI, as this is already the case for SISH.
III-C Pre-Trained Networks
Besides the deep networks published as part of the search engine methods, we used two other pre-trained networks as the backbone, as needed. One is DenseNet121 [55] and the second one is KimiaNet [51].
III-D Datasets
We conducted our benchmarking on four internal and three public datasets. The internal datasets are collected from the Mayo Clinic each from a different organ with different labels: breast epithelial tumor subtypes, fatty liver disease, cutaneous squamous cell carcinoma (cSCC) subtypes, and colorectal polyps. Table II gives a detailed overview of the number of patients, WSIs, and classes in these public/internal datasets.
TABLE II: Validation Datasets. The number of patients, WSIs, and classes in each dataset used for benchmarking the search engines.
Internal Dataset Number of Patients Number of WSIs Number of Classes
Breast Epithelial Tumors 74 74 16 + normal
Fatty Liver Disease 326 326 2 + normal
cSCC 660 660 3 + normal
Colorectal Polyps 209 209 3
Public CAM16 [56] 129 129 2
BRACS [57] 87 87 3
PANDA [58] 1000 1000 6
  • cSCC = cutaneous squamous cell carcinoma
III-D1 Internal Datasets
III-D2 Public Datasets
Working on rather small datasets with a limited number of classes/labels may not fully replicate the real-world application of image search if applied on larger WSI repositories. However, large WSI collections on high-performance storage devices with access to GPU computation does not seem to be available yet.
IV Analysis and Results
We first conduct analysis of the search engines in terms of their algorithmic structure, search capabilities, training, and test data used. Subsequently, we conducted extensive experiments to compare their search performance.
@@ IV Analysis and Results
Analysis – As for algorithmic structure comparison, we compared the patching methods they used to break down the WSI and perform indexing and matching. Overall, by comparing their search capabilities, we determined whether they are able to process/search for both patch and WSI matching. Training and testing data comparisons encompass a detailed inventory of the data each search engine was trained with or validated on.
Experiments – We used both high-quality clinical datasets and well-known public datasets to quantify the search performance of each method. We examined accuracy, time measurements for indexing and search, robustness, and storage requirements as main performance benchmarking. We standardized the implementation by utilizing a consistent computational machine and refrained from employing multiprocessing techniques, ensuring equitable time readings for all methods.
IV-A Analysis of Algorithmic Structure
@@
Table III offers an overview of the algorithmic structures employed by the four search frameworks (also see Figures 1,2,3 and 4 that visualize only the first three columns of Table III). Each approach is detailed across six key stages: Divide, Features, Encoding, Space, Matching, Speed, and Post-Processing. To analyze the algorithmic structure, we considered both the descriptions in the corresponding paper, as well as details of their implementation. We did not list “pre-processing” as all four methods start with segmenting the slide into tissue and non-tissue regions.
TABLE III: Algorithmic comparisons of all search methods (also see Figures 1,2, 3 and 4)
@@
Divide – Since using all WSI patches for the downstream tasks is inefficient in terms of both computation and storage (as this is the case for SMILY [15]), search methods must employ techniques to select a subset of the patches, also known as the “divide” stage in a divide & conquer attempt to master the gigapixel nature of WSIs [59]. While BoVW uses a random sampling approach to select visual words (which are much smaller than patches), Yottixel uses its “mosaic” approach, an unsupervised approach to get a small set of patches. SISH does not offer any new solution and uses Yottixel’s mosaic, the same as RetCCL. Yottixel performs color and spatial clustering to select representative patches, whereas RetCCL uses deep feature and spatial clustering, a much more expensive variation.
Features – Every search method has to utilize deep neural networks to obtain deep features from the image data, i.e., patches. RetCCL uses its clustering-guided contrastive learning (CCL) trained network using TCGA images to extract features for all patches prior to the mosaic generation step since it utilizes them for the divide step. In contrast, Yottixel and SISH both generate feature vectors only for the mosaic patches. Yottixel uses a pre-trained DenseNet, which can be swapped by any other network. SISH also extracts features using DenseNet, with the addition of an autoencoder trained on TCGA images. The rationale behind integrating the autoencoder alongside DenseNet in SISH is not sufficiently justified. It has been argued that SISH does not use any self-supervision for training the autoencoder, hence the paper may have introduced a misnomer [60]. Originally, BoVW used local binary patterns (LBP) as features but for the sake of comparison, we replaced it with an autoencoder so all the search engines utilize deep features.
@@
We should bear in mind that the Big-Oh notation is a theoretical upper bound with some practical limitations. For instance, comparing $N$ patches with $M$ other patches is bounded by $\mathcal{O}(N^{2})$ for $N\!>\!M$. However, if we use binary features (like Yottixel’s barcodes) and measure similarity via Hamming distance, then the actual time is rather bounded by the $\mathcal{\theta}(N^{\alpha})$ with $\alpha\in(0,1]$.
Post-Processing – Generally, the search process concludes once results have been retrieved and sorted based on their similarity to the query. However, both SISH and RetCCL utilize an additional ranking algorithm after the search, which eliminates patches deemed less likely to be similar to the query. The intuition behind ranking is to find the most promising patches coming from the search engine on the basis of the uncertainty [3]. It relies on three helper functions Weighted Uncertainty (the uncertainty for each patch by computing their entropy), a Clean function (to remove outliers and the patches that are less similar to the query) and a Filtered-By-Prediction function (to remove patches with low likelihoods).
The inclusion of ranking as a post-processing step is evidently motivated by a lack of confidence in the search results. This approach seems misguided, as relying on post-search ranking implies a lack of trust in the initial results. The authors of SISH stress the significance of ranking for their outcomes, indicating that incorporating an autoencoder into the Yottixel pipeline has not yielded any discernible benefits, despite the paper emphasizing the autoencoder’s role.
IV-B Analysis of Search Capabilities
@@
It is an indispensable capability of a search engine in histopathology to be able to perform both patch search and WSI search. WSI-to-WSI matching compares representative WSI patches to representative WSI patches in absence of ROI delineations. These representative WSI matches, like Yottixel’s mosaic, are expected to be the result of some unsupervised patch selection method. The query, coming from the pathologist, could be the entire WSI or selected ROI but the indexed WSIs are generally not expected to have ROIs, especially for large repositories. Comprehensive WSI-to-WSI comparison is important as several malignancies, or multiple staging/grading quantities may be present in the image (see the Fig. S1 in Supplementary Materials that illustrates a simple case of prostate Gleason Scoring that may be falsified if we remove some of representative WSI patches).
Clearly, WSI-2-WSI matching (WSI search) is a more difficult task compared to patch-to-patch matching. Also, WSI search is pivotal for patient tissue comparison. Both BoVW and Yottixel stand out as comprehensive solutions with the ability to perform both patch and WSI searches. In contrast, SISH and RetCCL, while being able to process and search patches due to using the Yottixel’s mosaic, lack the capability to perform WSI-2-WSI matching. This in fact stems from their dependence on ranking patches after the search to find the label for the query WSI; ranking may separate the mosaic patches (some of them get high and some other low ranks). Therefore, although the label found by SISH and RetCCL may be correct based on top-$n$ patches, this does not constitute true WSI-2-WSI matching as the decision is not based on all mosaic patches. In the case of SISH, even without post-search ranking, the tree used for fast processing cannot be restricted to top-$n$ WSIs.
In summary, BoVW creates a histogram for an image (either a patch or a WSI). The histogram can be calculated for both patch and WSI (Figure 9, left). Hence, BoVW can process and match both patches and WSIs. Yottixel creates a mosaic of patches for each WSI and uses the “median of minimum distances” to get a single distance metric for WSI comparison (Figure 9, middle); Hence, Yottixel can match both patches and WSIs. Both SISH and RetCCL use Yottixel’s mosaic but their ranking algorithms essentially force them to perform patch matching although the ranking may be configured to select one patch per WSI in order to emulate WSI selection. They take a query patch and compare it with every other patch in every other WSI (Figures 9, right). SISH pre-raking tree processing is not WSI-conform either. Hence, SISH and RetCCL are rather patch classifiers when using post-search ranking, hence not usable for patient tissue comparison in a clinical setting.
@@
IV-C Analysis of Training/Test Data
The data used for any type of training, the test data, and the user feedback experiments are all important factors in evaluating search engines. Table IV offers an overview of the data utilized by the four search engines. BoVW uses Kimia Path24555https://github.com/RhazesLab/kimia_path24 as testing data, comprising 24 different anatomic sites [64]. Notably, BoVW, as reported in [1], does not use any deep features and does not conduct user tests. Yottixel uses a pre-trained model, namely DenseNet. Hence, it has been validated on entire TCGA images, both diagnostic slides (FFPE slides) and frozen sections, in addition to some private data from UPMC, accompanied by some user tests involving three pathologists. The data Yottixle has been validated on covers a wide range of anatomic sites, totaling 27 primary sites, and an extensive number of subtypes. It is notable that since the pre-trained DenseNet121 has been trained on ImageNet, which is a non-pathological dataset, all Yottixel tests can be considered external validation. SISH draws its training data from TCGA and conducts testing with a diverse dataset comprising TCGA, CPTAC, and BWH data, without involving user tests. It focuses on a reduced set of anatomic sites (13) but accounts for 56 different subtypes. It is not clear how the training and testing TCGA data for SISH have been separated. RetCCL uses TCGA and PAIP data for training and only TCGA data for testing, without user tests. RetCCL considers 13 anatomic sites but focuses on a more limited set of subtypes.
TABLE IV: Comparison of Training and Test Data that each Search Engine has reported in corresponding papers.
Training Data Test Data Users